OpenStudy (anonymous):

Convert the improper fraction to a mixed number and reduce to lowest terms. 60 7 = Convert an improper fraction to a mixed number with three unknowns 7199 4.F.4.5.4

OpenStudy (anonymous):

what is the answer....... i dont get it sorry

OpenStudy (anonymous):

i don't quite understand your question but from what i understand. your objective is basically to reduce and turn an improper fraction to a proper fraction. how you got about that is: 60/7 = ? how many times can 7 go into 60? well, it goes 8 times because [7*8 = 56] so, 8 is the whole number part, then we need to find the fractional remainder. 60-56 = 4. and, the problem itself is over 7, so keep the denominator as 7. so your answer is: 8 and 4/7. a method to check your answer: in calculator type in 60/7 which should give you the answer as: 8.5714285714285714285714285714286 after that, take the remainder part and type in: 4/7 the answer you should get is: 0.57142857142857142857142857142857 now add 8 to it. thus, your answer is: 8.5714285714285714285714285714286 so in terms of decimals we are right.

OpenStudy (anonymous):

good Question. :)